# Repository-level Context Code Translation Benchmark Targeting Rust

### v5

@@ 3.3. Resulting Benchmark
Repository-level Dependency. RustRepoTrans focuses on code translation tasks with rich repository context, setting it apart from previous benchmarks. Unlike artificially constructed data or data from Q&A websites, real project data exhibits more complex dependency relationships, including function, data type, and variable dependencies, which are absent in existing datasets as shwon in Table 1, revealing that only RustRepoTrans includes these crucial elements. This inclusion makes RustRepoTrans a more realistic benchmark, suitable for evaluating LLMs that must account for intricate file-level interactions and contextual dependencies.
Rust Programming in Realistic Scenarios. RustRepoTrans and CodeTransOcean are the only benchmarks specifically targeting code translation to Rust, as listed in Table 1. However, RustRepoTrans is derived from GitHub projects, making it more reflective of real-world development than CodeTransOcean, which relies on data from programming competition websites.
4. Evaluation
Based on RustRepoTrans, we further investigate the performance of the state-of-the-art LLMs on repo-level code translation task for Rust. Specifically, we focus on the following RQs.
  • RQ1 (LLMs Performance): How do the state-of-the-art LLMs perform on RustRepoTrans in terms of translation effectiveness?
@@ 4. Evaluation
  • RQ3 (Difficulty Analysis): What factors affect the difficulty of the translation tasks?
  • RQ4 (Failure Analysis): What types of errors do LLMs encounter on RustRepoTrans, and what factors contribute to these translation failures?
  • RQ5 (Key Capabilities): Beyond translation accuracy, what essential capabilities do LLMs demonstrate on RustRepoTrans, including noise robustness, syntactical difference identification, and code simplicity?
4.1. Experimental Setup
Model Selection. We selected four state-of-the-art LLMs that have been widely studied in recent code translation research or were recently released and have demonstrated strong performance on benchmark leaderboards (e.g.,SuperCLUE (Xu et al., 2023)). Table 2 provides an overview of the four LLMs used in our experiments, along with their release dates (Column “Time”) and model sizes (Column “Size”). As shown in Table 2, our study includes both open-source and closed-source LLMs, as well as models designed for general-purpose and code-specific applications. Due to resource constraints, we selected open-source models with sizes below 20B parameters. Note that our primary contribution is a new benchmark designed to provide a more challenging and realistic evaluation of LLMs’ code translation capabilities, closely simulating real-world development scenarios. Expanding the scope to include a larger and more diverse set of models is reserved for future work.
Implementation Details. For open-source models, we obtained and executed the released versions from their official repositories based on provided documentation with greedy sampling (Chen et al., 2015) as our generation strategy. All evaluations were conducted on an NVIDIA A800 80GB GPU. For closed-source LLMs, we accessed each model through its respective API interface (OpenAI, 2024; Claude, 2024) as of September 2024. To achieve results similar to greedy decoding, we set the “temperature” hyperparameter to 00.
4.2. RQ1: LLMs Performance
We evaluated four selected LLMs on RustRepoTrans to assess their translation accuracy and error-correction abilities for complex, repository-level code translation tasks targeting Rust.
4.2.1. Design
@@
Evaluation Process. Each selected LLM was tasked with translating code into Rust, specifically focusing on 375 tasks. For each translation task, a corresponding set of test cases was employed to evaluate the correctness of the generated outputs. To ensure a fair comparison, the same prompt was used for each model, carefully designed based on established best practices in code translation tasks, as illustrated in Fig. 3. These best practices have been shown to enhance translation accuracy in similar studies (Macedo et al., 2024b; Khan et al., 2024b; Pan et al., 2024; Yang et al., 2024). The prompt includes elements such as Instruction for translation and Translation’s required information(including source code, target function signature and target function dependencies)
Table 2. Studied LLMs
Model Type Model Company
Open-source
Time Size
General LLM Claude-3.5 (Anthropic, 2024) Anthropic 2024.6 -
GPT-4 (OpenAI, 2023) OpenAI 2023.6 -
Llama-3.1-8B (Meta, 2024) Meta 2024.7 8B
Code LLM DeepSeekCoderV2-16B (DeepSeek, 2024) DeepSeek 2024.6 16B
@@
Given that existing LLM-based code translation research often incorporates feedback on translation error messages to enhance performance (Yang et al., 2024; Pan et al., 2024), further experiments were conducted to understand the extent to which errors in the generated code samples can be corrected by LLMs with feedback.
It is important to note that the goal is to evaluate the LLMs’ intrinsic ability to self-correct translation errors, as well as to explore how many translation errors in RustRepoTrans can be automatically corrected by LLMs. Therefore, in addition to self-debugging, the best-performing model on RustRepoTrans, Claude-3.5, was also used to debug the translation outputs of all models. The debugging prompt (Prompt-Fix) shown in Fig. 3 includes Instruction for previous translation, Instruction for debugging, Incorrect translation and error details and Translation’s required information, which are adapted from previous work (Pan et al., 2024; Yang et al., 2024).
Metrics. Two key metrics were utilized to measure performance:
  • Pass@k: This widely-used metric (Chen et al., 2021) calculates the percentage of tasks correctly solved based on $k$ generated code samples per task, as expressed in Equation 1. A task is considered solved if at least one of the generated code samples passes all the corresponding test cases. Following recent research (Pan et al., 2024), the focus was on calculating the Pass@1 metric, where $k=1$. This approach reflects the model’s ability to produce a correct translation on the first attempt.
[…]
  • DSR@k: The DSR@k (Debugging Success Rate@k) metric, proposed in previous work (Yan et al., 2023), evaluates whether the generated code successfully executes and produces the expected results (i.e.,passing all test cases) within $k$ rounds of debugging, as expressed in Equation 2. $S(i,k)=1$ if the $i^{th}$ code sample succeeds within $k$ attempts; otherwise $S(i,k)=0$. In this study, DSR@k metrics were calculated with $k=1$, allowing for one debugging attempt.
[…]
Figure 3. Translating Prompt and Debugging Prompt
4.2.2. Results
@@
Fig. 4 presents the Pass@1 and DSR@1 performance of each LLM on RustRepoTrans. This benchmark challenges models with tasks that include complex dependencies, offering insights into each model’s initial translation accuracy and post-debugging improvement potential.
@@
Initial Translation Performance (Pass@1). In the initial code translation task, Claude-3.5 substantially outperformed the other LLMs, achieving 14.2% higher accuracy than GPT-4, 26.7% higher than DeepSeekCoderV2-16B, and 30.2% higher than Llama-3.1-8B on the Pass@1 metric. This indicates Claude-3.5’s comparative strength in handling repository-level code translation to Rust. However, even Claude-3.5 achieved only a 45.3% Pass@1 score, emphasizing the difficulty of RustRepoTrans and the challenges inherent in translating code with complex dependencies.
Post-debugging Performance. After a single round of self-debugging, each model’s DSR@1 score showed significant improvement, confirming previous research that LLMs can effectively leverage compiler feedback for code translation accuracy (Yan et al., 2023; Yang et al., 2024). However, improvements from debugging by Claude-3.5 surpassed those from self-debugging, with Llama-3.1-8B showing only a 5.6% increase from self-debugging compared to a 24.8% improvement from Claude-3.5. This underscores Claude-3.5’s superior debugging capabilities in complex code translation.
@@
Post-debugging performance rankings revealed Claude-3.5 leading at 56.5%, followed by GPT-4 at 44.5%. Notably, Llama-3.1-8B (38.1%) surpassed DeepSeekCoderV2-16B (36.0%), indicating its initial errors were more correctable. Despite Claude-3.5 ’s strong performance, 44.5% of its cases remained unresolved, highlighting the inherent challenges of self-correction in intricate translation tasks. In terms of relative improvement, Llama-3.1-8B exhibited the highest percentage gain at 186%, indicating its initial errors were highly correctable. Conversely, Claude-3.5 showed the smallest improvement at 13.0%, suggesting its initial outputs were closer to correct execution. Overall, even after self-debugging, Claude-3.5 maintained its leading position, emphasizing its robust capabilities in repository-level code translation to Rust.
Figure 4. Pass@1, DSR@1(self-debugging) and DSR@1(debug by Claude-3.5) on RustRepoTrans
4.2.3. Summary.
@@
The results reveal that RustRepoTrans effectively challenges LLMs in complex code translation, as even the best model, Claude-3.5, only achieves a Pass@1 of 45.3%, and DSR@1 of 56.5%, highlighting the limitations of LLMs.
4.3. RQ2: RustRepoTrans Effectiveness
@@
We evaluated performance on RustRepoTrans and a prior Rust translation benchmark without repository-level context, demonstrating RustRepoTrans’s effectiveness in assessing models for complex, real-world translation scenarios. Our analysis is based on CodeTransOcean (Yan et al., 2023), which includes translation tasks across 45 programming languages, including Rust, without repository-level dependencies. To ensure a fair and efficient comparison, we constructed a subset of CodeTransOcean by selecting translation pairs where Rust is the target language and C, Java, or Python is the source language, removing those with non-compilable Rust code. From the remaining pairs, we randomly sampled 300 (100 per language) to match our dataset’s scale for comparative experiments.
Claude-3.5 was selected for its top performance in RustRepoTrans. It generated one translation per task, evaluated using corresponding tests. We computed the average Pass@1 across source languages and compared it with RQ1. Prompts followed those in Fig. 3 but excluded dependency-related instructions. Hyperparameters were the same as in RQ1, with the temperature set to zero.
4.3.2. Results
@@
The results presented in Table 3 reveal significant differences in the Pass@1 scores of Claude-3.5 across various datasets. Notably, while Claude-3.5 achieves high accuracy rates—ranging from 67.0% to 80.0%—on established benchmarks, its performance drastically drops when evaluated on RustRepoTrans.
@@
Specifically, the Pass@1 scores for RustRepoTrans are significantly lower, with rates of 48.3% for C to Rust, 39.3% for Java to Rust, and 41.7% for Python to Rust translations, averaging 42.5%. This indicates a decline of approximately 25.3% to 36.7% compared to the higher scores observed on CodeTransOcean. This notable decrease highlights the increased complexity and unique challenges posed by RustRepoTrans, reinforcing the need for models to effectively manage repository-level dependencies and navigate intricate code structures in real-world.
Table 3. Pass@1 Comparison of Claude-3.5 in C, Java, and Python to Rust on CodeTransOcean and RustRepoTrans
Source Language Dataset Pass@1(Claude-3.5)
C CodeTransOcean 80.0%
RustRepoTrans 48.3%(↓31.7%)
Java CodeTransOcean 76.0%
RustRepoTrans 39.3%(↓36.7%)
Python CodeTransOcean 67.0%
RustRepoTrans 41.7%(↓25.3%)
4.3.3. Summary
@@
Our benchmark’s repository-level context poses greater challenges for LLMs, reducing Claude-3.5’s Pass@1 from 74.3% to 43.5% (a 30.8% drop) in average. This underscores the benchmark’s effectiveness in evaluating real-world code translation and the need for models to better handle project dependencies.
4.4. RQ3: Difficulty Analysis
To understand the challenges that RustRepoTrans presents to LLMs and how they contribute to performance drops, we analyze the code translation results of LLMs across different tasks.
4.4.1. Design
We analyzed both successfully and unsuccessfully translated tasks using Claude’s results, chosen for its superior performance. Our goal was to identify the types of tasks where it succeeds or fails. Based on our dataset characteristics, we focused on three key factors that influence translation difficulty: the number of dependencies, and the lines of code in both the source and target languages. We conducted statistical analyses of these factors to assess their impact on translation success.
4.4.2. Results
Fig. 5 illustrates the distribution of successful and unsuccessful tasks with Claude based on dependencies and source/target code lines. Successful translations are concentrated on the left side, indicating fewer dependencies and shorter code lines. In contrast, unsuccessfully translated pairs show a wider distribution. Specifically, failed translations have, on average, 138.7% more dependencies, 134.6% more lines of code in the target language, and 108.4% more lines of code in the source language compared to successful ones. This suggests that failed function pairs present more challenging translation tasks, resulting in lower success rates.
Figure 5. Success and Failure Distribution of Claude by Dependencies and Source/Target Code Lines
4.4.3. Conclusion
Higher dependencies and longer code lines challenge LLMs, resulting in lower translation success rates.
4.5. RQ4: Failure Analysis
@@ 4.5.1. Design
Automatic Error Outcome Analysis. We collected all unsuccessful translations from LLMs in RQ1 (Section 4.2), resulting in 1,114 code samples that failed to pass all test cases. To identify where LLMs struggle in translating code to Rust, we conducted an automated analysis following the methodology of Pan et al. (Pan et al., 2024), categorizing the unsuccessful translations into four error outcomes: compilation errors (code fails to compile), runtime errors (code compiles but encounters exceptions), functional errors (code executes successfully but fails test cases), and non-terminating execution (code runs indefinitely or waits for input).
@@
Manual Error Causes Analysis. We further conducted a manual analysis of the error causes for the unsuccessful translations with compilation errors (1,056 code samples), which accounted for the vast majority, 94.8%, of the failures as shown in Figure 4 (see Section 4.5.2). Using an open coding approach (error codes, 2024), we examined the error messages for each failing result and classified each compilation error into specific categories. If an error could not fit into an existing category, we created a new one or adjusted the definition of an existing category to include it, followed by revising and reannotating all relevant samples.
Notably, the same type of error could belong to different categories depending on the context. This iterative annotation process involved collaborative discussions to refine the categories and ensure consistent classification. We regularly reviewed and reorganized the categories to maintain clarity in our classification system.
4.5.2. Results
@@
Error Outcome Distribution. Table 4 shows the error outcome distribution from the RQ1 experiment on RustRepoTrans, alongside data from Pan et al. (Pan et al., 2024) for translations to other programming languages (C, C++, Python, Java).
@@
In RustRepoTrans, compilation errors account for 94.8% of failures, significantly higher than the 58.3% to 83.3% observed in other benchmarks. This emphasizes the challenge LLMs face when translating to Rust, a language known for stringent compile-time checks due to its ownership model and borrow checker. Unlike other benchmarks, RustRepoTrans recorded no runtime or non-terminating errors, indicating that Rust translations primarily fail at compile-time, underscoring its strong emphasis on memory safety.
@@
The number of compilation errors in unsuccessful translations ranged from 1 to 193, with an average of 9.1 and a median of 4, which complicates debugging. In contrast to languages like Python, where the compiler stops at the first error, the Rust compiler reports all errors, making issue resolution more challenging. We categorized the types of compilation errors based on error codes from (error codes, 2024). Table 5 presents the top 10 compilation errors encountered, with the five most common errors occurring over 1,000 times each. Many of these stem from using non-existent, unimplemented, unresolved, or undeclared elements, highlighting the hallucination phenomenon also seen in other repository-level code generation tasks (Luo et al., 2024; Wang et al., 2024b; Du et al., 2023; Wang et al., 2024a; Liu et al., 2024).
@@
Table 4. The comparison of the proportions of error types between RustRepoTrans and (Pan et al., 2024).CE: Compilation Errors, RE: Runtime Errors, FE: Functional Errors, NTE: Non-terminating Execution
@@
RustRepoTrans C, Java, Python Rust 94.8% 0% 5.2% 0%
Table 5. Top 10 Compilation Errors Reported by the Rust Compiler on RustRepoTrans
@@
E0432 3,504 An import was unresolved.
E0277 2,540
@@
E0433 1,345 An undeclared crate, module, or type was used.
E0609 796 Attempted to access a nonexistent field in a struct.
E0308 744 Expected type did not match the received type.
E0061 222
@@
E0412 174 A used type name is not in scope.
E0282 79
The compiler could not infer a type and asked for
a type annotation.
Error Cause Taxonomy. Fig. 6 illustrates the ten error cause categories we identified for failed code translations on RustRepoTrans, specifically those with compilation errors. We further classified these into three main types:
@@
  • Failing to Understand Differences Between Languages. This type relates to the LLM’s inadequate comprehension of the distinctions between the source and target programming languages. It covers issues such as syntax variations, differences in functions, variables, data types, and import paths.
@@
Figure 6. Error Cause Taxonomy for Failed Code Translations on RustRepoTrans in LLMs
@@
Data Type Misinterpretation. This error cause reflects the LLM’s limitations in type inference, particularly evident in Rust’s strict type system. Unlike Python, where variables can change types freely, or C, which allows implicit conversions, Rust requires exact type matches, making translation errors more prominent. LLMs may attempt incompatible assignments, perform arithmetic on mismatched types, or incorrectly treat non-collection variables (like integers) as if they were indexable. In the example below, the model interprets imap_connected_here as a boolean, causing a type mismatch that Rust’s compiler disallows. This illustrates how Rust’s strict typing exposes LLMs’ type inference limitations, often masked in more permissive languages like Python.
@@
Variable State Misinterpretation. This error cause stems from the LLM’s misunderstanding of variable states, leading to several misinterpretations. Common issues include treating undeclared variables as declared, making simultaneous mutable borrows, using the same variable as both mutable and immutable, and incorrectly treating private variables as public. In the example below, the model attempts to pass h as both an immutable reference (&h) and a mutable reference (&mut h), which Rust’s strict borrowing rules prohibit, resulting in a compilation error. These instances highlight the LLM’s limitations in tracking variable states and usage contexts, often leading to significant translation errors.
fn hashit(sha: usize, n: usize, id: &[u8], w: &mut [u8]) -> bool    # hashit function definition hashit(sha, date, &h, &mut h);   # error: using h as both mutable and immutable parameter
Context Misinterpretation. This error cause arises from the LLM’s misunderstanding of context. It includes issues such as failing to recognize when code is within a struct’s built-in function, using asynchronous operations in a synchronous block, or attempting modifications inside a function without the mutable modifier. In the example below, the model incorrectly assumes that batch is a built-in function of a struct. The error occurs when it tries to access self.left(batch.clone()) without using the keyword self, which is necessary in this context.
@@
Syntactic Differences Misinterpretation. This error cause relates to the LLM’s understanding of syntactic differences between programming languages. It occurs when the model incorrectly assumes that syntax from the source language is valid in the target language. Examples include using chained assignments (allowed in Python, C, and Java) in Rust, attempting to use the goto keyword from C, or employing an unsupported ternary operator ‘‘condition ? expr1 : expr2’’ in Rust.
Function Differences Misinterpretation. This error cause pertains to the LLM’s understanding of functional dependency differences between programming languages. It arises when the model incorrectly assumes that callable functions in the target language are identical to those in the source language. Common issues include assuming a function present in the source language also exists in the target language, that it shares the same scope, or that the invocation method is identical. In the example below, the first line results in an error because the LLM fails to recognize that get should be called within the stock_str namespace.
ret.text = get(context, StockMessage::NoMessages)?;       # error ret.text = stock_str::get(context, StockMessage::NoMessages)?;       # correct

### v6

@@ III-C Resulting Benchmark
Repository-level Dependency. RustRepoTrans focuses on code translation tasks with rich repository context, setting it apart from previous benchmarks. Unlike artificially constructed data or data from Q&A websites, real project data exhibits more complex dependency relationships, including function, data type, and variable dependencies, which are absent in existing datasets as shwon in Table I, revealing that only RustRepoTrans includes these crucial elements. This inclusion makes RustRepoTrans a more realistic benchmark, suitable for evaluating LLMs that must account for intricate file-level interactions and contextual dependencies.
Rust Programming in Realistic Scenarios. RustRepoTrans and CodeTransOcean are the only benchmarks specifically targeting code translation to Rust, as listed in Table I. However, RustRepoTrans is derived from GitHub projects, making it more reflective of real-world development than CodeTransOcean, which relies on data from programming competition websites.
IV Evaluation
Based on RustRepoTrans, we further investigate the performance of studied LLMs on repo-level code translation task for Rust. Specifically, we focus on the following RQs.
  • RQ1 (LLMs Performance): How do the studied LLMs perform on RustRepoTrans in terms of translation effectiveness?
@@ IV Evaluation
  • RQ3 (Failure Analysis): What types of errors do LLMs encounter on RustRepoTrans, and what factors contribute to these translation failures?
  • RQ4 (Key Capabilities): Beyond translation accuracy, what essential capabilities do LLMs demonstrate on RustRepoTrans, including noise robustness, syntactical difference identification, and code simplicity?
IV-A Experimental Setup
Model Selection. As shown in Table II, we selected seven presentative LLMs that have been widely studied in recent research including general LLM and code LLM, open-source and closed-source models, as well as reasoning models and non-reasoning models to conduct a comprehensive evaluation of LLM’s code translation capabilities in real-world scenarios.
Implementation Details. Due to resource constraints, for open-source models with sizes below 20B parameters, we obtained and executed the released versions from their official repositories with greedy sampling [44] as our generation strategy. All evaluations were conducted on an NVIDIA A800 80GB GPU. For closed-source LLMs and other open-source models, we accessed each model through their official API interface [45, 46, 47, 48, 49]. To achieve results similar to greedy decoding, we set the “temperature” hyperparameter to 0.
IV-B RQ1: LLMs Performance
We evaluated studied LLMs on RustRepoTrans to assess their translation accuracy and self-debugging abilities for repository-level context code translation tasks targeting Rust.
IV-B1 Design
@@
Evaluation Process. Each selected LLM was tasked with translating code into Rust, specifically focusing on 375 tasks. For each translation task, a corresponding set of test cases was employed to evaluate the correctness of the generated outputs. To ensure a fair comparison, the same prompt was used for each model, carefully designed based on established best practices in code translation tasks, as illustrated in Fig. 4. These best practices have been shown to enhance translation accuracy in similar studies [50, 51, 16, 35]. The prompt includes elements such as instruction for translation and translation’s required information(including source code, target function signature and target function dependencies)
TABLE II: Studied LLMs
Model Type Model Name Open-source reasoning Time Size
General LLM DeepSeek-R1-0528 [47] 2025.5 671B
DeepSeek-V3-0324 [48] 2025.3 671B
Claude-3.5-Sonnet[52] 2024.6 -
GPT-4 [53] 2023.6 -
Llama-3.1-8B [54] 2024.7 8B
Code LLM Qwen-2.5-coder-32B [55] 2024.9 32B
DeepSeekCoderV2-16B [56] 2024.6 16B
@@
Given that existing LLM-based code translation research often incorporates feedback on translation error messages to enhance performance [35, 16], further experiments were conducted to understand the extent to which errors in the generated code samples can be corrected by LLMs with feedback. The debugging prompt (Prompt-Fix) in Fig. 4 includes instruction for previous translation, instruction for debugging, incorrect translation and error details and translation’s required information, adapted from previous work [16, 35].
Metrics. In line with previous work [16, 23], two key metrics: Pass@1 and DSR@1 were utilized. Pass@1 [57] reflects the LLMs’ ability to produce correct translation on the first attempt, while DSR@1 [23] reflects the model’s ability to produce correct translation allowing for one debugging attempt.
Figure 4: Translating Prompt and Debugging Prompt
Figure 5: Pass@1, DSR@1(self-debugging) on RustRepoTrans
IV-B2 Results
@@
Fig. 5 presents the Pass@1 and DSR@1 performance of each LLM on RustRepoTrans. This benchmark challenges models with tasks that include complex dependencies, offering insights into each model’s initial translation accuracy and self-debugging improvement potential.
@@
Initial Translation Performance (Pass@1). In initial code translation task, DeepSeek-R1 substantially outperformed other LLMs, achieving 51.5% accuracy. Among non-reasoning LLMs, DeepSeek-V3 achieving highest accuracy, only 1.4% lower than DeepSeek-R1 and 6.6% higher than Claude-3.5, 15.7% higher than Qwen-2.5-coder-32B, 20.8% higher than GPT-4, 33.3% higher than DeepSeekCoderV2-16B, and 36.8% higher than Llama-3.1-8B on the Pass@1 metric. However, even DeepSeek-R1 achieved only a 51.5% Pass@1 score, emphasizing the difficulty of RustRepoTrans and the challenges inherent in translating code with complex dependencies.
@@
Self-debugging Performance. After a single round of self-debugging, each model’s DSR@1 score showed significant improvement, confirming previous research that LLMs can effectively leverage compiler feedback for code translation accuracy [23, 35]. The self-debugging performance rankings revealed DeepSeek-R1 leading at 62.1%, followed by DeepSeek-V3 at 58.7%. Despite the strong performance of DeepSeek-R1, 37.9% of its cases remained unresolved, highlighting the inherent challenges of self-correction in intricate translation tasks. In terms of relative improvement, Claude-3.5 exhibited the highest percentage gain at 29.9%, indicating its strong capabilities of self-debugging in repository-level context code translation. Overall, even after self-debugging, DeepSeek-R1 maintained its leading position, emphasizing its robust capabilities in incremental code translation scenario to Rust.
IV-B3 Summary
@@
Results reveal that RustRepoTrans effectively challenges LLMs in complex code translation, as even the best model, DeepSeek-R1, only achieves a Pass@1 of 51.5%, and DSR@1 of 62.1%, highlighting the limitations of LLMs.
IV-C RQ2: RustRepoTrans Effectiveness
@@ IV-C1 Design
We evaluated performance on RustRepoTrans and a prior Rust translation benchmark without repository-level context, demonstrating RustRepoTrans’s effectiveness in assessing models for complex, real-world translation scenarios. Our analysis is based on CodeTransOcean [23], which includes translation tasks across 45 programming languages, including Rust, without repository-level dependencies. To ensure a fair and efficient comparison, we constructed a subset of CodeTransOcean by selecting translation pairs where Rust is the target language and C, Java, or Python is the source language, removing those with non-compilable Rust code. From the remaining pairs, we randomly sampled 300 (100 per language) to match our dataset’s scale for comparative experiments.
We comprehensively evaluated top-performing models in RustRepoTrans (DeepSeek-R1, DeepSeek-V3, Claude-3.5, Qwen-2.5-coder-32B) by comparing their average Pass@1 and DSR@1 on a CodeTransOcean subset with RQ1 results. The prompts (Fig. 4) excluded dependency instructions, and the hyperparameters remained consistent with RQ1.
IV-C2 Results
@@
The results presented in Table III reveal significant differences in the Pass@1 and DSR@1 scores of LLMs across various datasets. Notably, while LLMs achieve high accuracy rates—with DeepSeek-R1 being the top performer, attaining a Pass@1 of 73.7% and a DSR@1 of 89.0%—on established benchmarks, their performance drastically drops when evaluated on RustRepoTrans.
@@
Specifically, the Pass@1 scores for RustRepoTrans are significantly lower, with a decline of approximately 16.2% to 30.8% compared to the higher scores observed on CodeTransOcean. Moreover, after self-debugging, the accuracy gap further widens, with a decline of 26.3% to 38.4% in DSR@1. These notable decreases highlight the increased complexity and unique challenges posed by RustRepoTrans, reinforcing the need for models to effectively manage repository-level dependencies and navigate intricate code structures in real-world scenarios. Furthermore, these complexities and challenges are less easily corrected through simple error message feedback.
It is also important to note that Claude-3.5 and DeepSeek-R1, which exhibit a clear performance gap on RustRepoTrans, demonstrate very similar results on CodeTransOcean. This indicates that existing stand-alone, function-level datasets can only evaluate a model’s basic code translation capability and are insufficient for effectively distinguishing between models with stronger code abilities. Therefore, a more challenging benchmark like RustRepoTrans, which incorporates repository-level context and targets incremental translation scenarios, is necessary to provide a finer-grained and multifaceted evaluation of the models’ code translation capabilities.
TABLE III: Pass@1 and DSR@1 Comparison of LLMs’ performance on CodeTransOcean and RustRepoTrans
Model Dataset Pass@1 DSR@1
DeepSeek-R1 CodeTransOcean 73.7% 89.0%
RustRepoTrans 51.5%(↓22.2%) 62.1%(↓26.9%)
DeepSeek-V3 CodeTransOcean 66.3% 85.0%
RustRepoTrans 50.1%(↓16.2%) 58.7%(↓26.3%)
Claude-3.5 CodeTransOcean 74.3% 87.7%
RustRepoTrans 43.5%(↓30.8%) 56.5%(↓31.2%)
Qwen-2.5-coder-32B CodeTransOcean 56.7% 77.3%
RustRepoTrans 34.4%(↓22.3%) 38.9%(↓38.4%)
IV-C3 Summary
@@
Our benchmark’s repository-level context poses greater challenges for LLMs, reducing their Pass@1 and DSR@1 scores by 16.2% to 30.8% and 26.3% to 38.4%, respectively, while simultaneously providing greater differentiation between models that perform similarly on existing datasets. This underscores the benchmark’s effectiveness in evaluating real-world code translation capabilities and highlights the need for models to better handle project dependencies.
IV-D RQ3: Failure Analysis
@@ IV-D1 Design
Automatic Error Outcome Analysis. We collected all unsuccessful translations from LLMs in RQ1 (Section IV-B), resulting in 1,748 code samples that failed to pass all test cases. To identify where LLMs struggle in translating code to Rust, we conducted an automated analysis following the methodology of Pan et al. [16], categorizing the unsuccessful translations into four error outcomes: compilation errors (code fails to compile), runtime errors (code compiles but encounters exceptions), functional errors (code executes successfully but fails test cases), and non-terminating execution (code runs indefinitely or waits for input).
@@
Manual Error Causes Analysis. We further conducted a manual analysis of the error causes for the unsuccessful translations with compilation errors (1,614 code samples), which accounted for the vast majority, 92.3%, of the failures as shown in Table IV (see Section IV-D2). Using an open coding approach [58], we examined the error messages for each failing result and classified each compilation error into specific categories. If an error could not fit into an existing category, we created a new one or adjusted the definition of an existing category to include it, followed by revising and reannotating all relevant samples. Notably, the same type of error could belong to different categories depending on the context. This iterative annotation process involved collaborative discussions to refine the categories and ensure consistent classification. We regularly reviewed and reorganized the categories to maintain clarity in our classification system.
IV-D2 Results
@@
Error Outcome Distribution. Table IV shows the error outcome distribution from RQ1 on RustRepoTrans, alongside data from Pan et al. [16] for translations to other programming languages (C, C++, Python, Java).
@@
In RustRepoTrans, compilation errors account for 92.3% of failures, significantly higher than the 58.3% to 83.3% observed in other benchmarks. This emphasizes the challenge LLMs face when translating to Rust, a language known for stringent compile-time checks due to its ownership model and borrow checker. Unlike other benchmarks, RustRepoTrans recorded no runtime or non-terminating errors, indicating that Rust translations primarily fail at compile-time, underscoring its strong emphasis on memory safety.
@@
In unsuccessful translations, the number of compilation errors ranged from 1 to 193 (average: 7.7, median: 3), complicating debugging. In contrast to languages like Python, where the compiler stops at the first error, the Rust compiler reports all errors, making issue resolution more challenging. We categorized the types of compilation errors based on error codes from [58]. Table V presents the top 10 compilation errors encountered. Many of these stem from using non-existent, unimplemented, unresolved, or undeclared elements, highlighting the hallucination phenomenon also seen in other repository-level code generation tasks [59, 60].
@@
TABLE IV: Comparison of the proportions of error types between RustRepoTrans and [16].CE: Compilation Errors, RE: Runtime Errors, FE: Functional Errors, NTE: Non-terminating Execution
@@
RustRepoTrans C, Java, Python Rust 92.3% 0% 7.7% 0%
TABLE V: Top 10 Compilation Errors by Rust Compiler on RustRepoTrans
@@
E0425 2,095 An unresolved name was used.
E0308 1,035 Expected type did not match the received type.
E0277 976
@@
E0609 906 Attempted to access a nonexistent field in a struct.
E0433 639 An undeclared crate, module, or type was used.
E0061 343
@@
E0252 290
Two items of the same name cannot be imported without
rebinding one of the items under a new local name.
E0432 287 An import was unresolved.
E0616 187 Attempted to access a private field on a struct.
Error Cause Taxonomy. We conduct open coding [61] with an iterative consensus on error messages to annotate error categories. During each iteration, two annotators with over 6 years of programming experience independently labeled errors by fully considering the error description of the error message and the context of the failed code translation and then resolved disagreements through discussion. After three rounds, the categories stabilized, ensuring consistent and reliable results. We randomly selected 50 error cases and asked another author to independently label them. The Cohen’s kappa [62] reached 0.885, which indicates almost perfect agreement, representing the validity and reproducibility of the error categories obtained. Fig. 6 shows the ten final error cause categories for failed code translations with compilation errors in RustRepoTrans, further grouped into three main types:
@@
  • Failing to Understand Differences Between Languages. This type relates to the LLM’s inadequate comprehension of the distinctions between the source and target languages. It covers issues such as syntax variations, differences in functions, variables, data types, and import paths.
@@
Figure 6: Error Cause Taxonomy for Failed Code Translations on RustRepoTrans in LLMs
@@
Data Type Misinterpretation. This error cause reflects the LLM’s limitations in type inference, particularly evident in Rust’s strict type system. Unlike Python, where variables can change types freely, or C, which allows implicit conversions, Rust requires exact type matches, making translation errors more prominent. LLMs may attempt incompatible assignments, perform arithmetic on mismatched types, or incorrectly treat non-collection variables (like integers) as if they were indexable. In the example below, the model interprets imap_connected_here as a boolean, causing a type mismatch that Rust’s compiler disallows. This illustrates how Rust’s strict typing exposes LLM’s type inference limitations, often masked in more permissive languages like Python.
@@
Variable State Misinterpretation. This error cause stems from the LLMs’ misunderstanding of variable states, leading to several misinterpretations. Common issues include treating undeclared variables as declared, making simultaneous mutable borrows, using the same variable as both mutable and immutable, and incorrectly treating private variables as public. In the example below, the model attempts to pass h as both an immutable reference (&h) and a mutable reference (&mut h), which Rust’s strict borrowing rules prohibit, resulting in a compilation error. These instances highlight the LLM’s limitations in tracking variable states and usage contexts, often leading to significant translation errors.
Function Definition: fn hashit(sha: usize, n: usize, id: &[u8], w: &mut [u8]) -> bool # hashit Function Calling: hashit(sha, date, &h, &mut h); # error: using h as both mutable and immutable
Context Misinterpretation. This error cause arises from the LLM’s misunderstanding of context. It includes issues such as failing to recognize when code is within a struct’s built-in function, using asynchronous operations in a synchronous block, or attempting modifications inside a function without the mutable modifier. In the example below, the model incorrectly assumes that batch is a built-in function of a struct. The error occurs when it tries to access self.left(batch.clone()) with using the keyword self, which is unnecessary in this context.
@@
Syntactic Differences Misinterpretation. This error cause relates to the LLM’s understanding of syntactic differences between programming languages. It occurs when the model incorrectly assumes that syntax from the source language is valid in the target language. Examples include using chained assignments (allowed in Python, C, and Java), attempting to use the goto keyword from C, or employing an unsupported operator ‘‘condition ? expr1 : expr2’’ in Rust.
Function Differences Misinterpretation. This error cause pertains to the LLM’s understanding of functional dependency differences between programming languages. It arises when the model incorrectly assumes that callable functions in the target language are identical to those in the source language. Common issues include assuming a function present in the source language also exists in the target language with identical function calling way. In the example below, the first line results in an error because the LLM fails to recognize that get should be called within the Self prefix.
resort(&mut items_vec);   # error Self::resort(&mut items_vec);   # correct